OpenStudy (anonymous):

How do you simplify: (8a^-6)^-2/3

OpenStudy (anonymous):

1/4a^4

OpenStudy (anonymous):

thanks! :)

OpenStudy (anonymous):

No problem. Don't forget to give me a badge and fan me.

OpenStudy (anonymous):

I'm new to this, so I did become your fan but I don't know how to give you a badge.

OpenStudy (anonymous):

Lol, I just join this just today.

OpenStudy (anonymous):

Beside my name, does it say that you could give me one?

OpenStudy (anonymous):

haha really?? But no, I don't see anything

OpenStudy (anonymous):

Yeah, oh the irony! You could search it up on how to do it or you could put "Best Response"

OpenStudy (tkhunny):

@vzfreakz Please read the Code of Conduct and abide by its wisdom. (8a^-6)^-2/3 \(\left(\dfrac{8}{a^{6}}\right)^{-2/3} = \left(\dfrac{a^{6}}{2^{3}}\right)^{2/3} = \left(\dfrac{a^{2}}{2}\right)^{2} = \dfrac{a^{4}}{2^2} = \dfrac{a^{4}}{4}\)